Construct a curve then shade the region corresponding to the obtained area from The weights (in kg) of 11-STEM B follow a normal distribution and has a mean of 50 and a standard deviation of 5. How many students have weights greater than 57?

1
Expert's answer
2022-06-13T11:40:03-0400
"P(X>57)=1-P(Z\le\dfrac{57-50}{5})"

"=1-P(Z\le1.4)=0.0808"